# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v2

@@ 1.2. General Setting
For simplicity, the pair $(f_{+},f_{-})$ will be assumed to be independent of $\Theta$ henceforth.
@@
(9) $X^{(n)}\overset{d}{=}\Phi(X^{(n)}_{1},X^{(n)}_{2})$,
where $X^{(n)}_{1},X^{(n)}_{2}$ are i.i.d. copies of $X^{(n-1)}$, which are independent of $\Phi$. Throughout the paper, equality is only required to hold in law; questions of couplings or constructions of $\{X^{(n)}\}$ will not be treated.
@@ 1.4.1. Main Results
(15) $\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<\infty,$
@@
$\sigma=-\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|]\,du+\int_{0}^{+\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds.$
@@
(16) $\int_{0}^{+\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds,$
define also the constant $a\geq 0$ by
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (16) holds, $\sigma=0$, and $a>0$. For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
@@
where $(X^{(N,n-1)}_{1},X^{(N,n-1)}_{2})$ are i.i.d. copies of $X^{(N,n-1)}$ and $\Phi^{(N)}$ is given by (8) with bias parameter $p=p^{(N)}$ satisfying $p^{(N)}=1/2+\theta N^{-1}$ for some $\theta\in\mathbb{R}$.
@@ 1.4.2. Idea of the Proof
Consistency, on the other hand, refers to the behavior of the equation under rescaling. Specifically, Proposition 10 below shows that if $\Phi$ is a smooth CDF, rescaled according to $\Phi_{\delta}(x)=\Phi(\delta x)$, then
(23) $(\mathscr{L}\Phi_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}\Phi(x)|^{2}+\delta^{3}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+\cdots\quad\text{as}\,\,\delta\downarrow 0.$
@@
which motivates the choices $\alpha=2$ and $\alpha=3$ used above and the appearance of the PDE’s (19) and (22). The approach of [7] provides a framework for making such a formal argument rigorous.
1.5. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
The results of this work suggest interesting new questions away from the critical point $p=\frac{1}{2}$. In view of the asymptotic expansion (23) and the PDE’s obtained in Theorems 1 and 2, the equation (11) shares some similarities with the reaction-diffusion equation
@@ 1.5. Variants of the Fisher-KPP Equation when $p\neq\frac{1}{2}$
A number of works in the PDE literature considered the equation (24) in the case when $\sigma=0$ and $a>0$. Using an auxiliary ODE, Enguiça, Gavioli, and Sanchez [17] established the existence of traveling waves propagating at any large-enough speed, as in the classical Fisher-KPP equation. The work of Audrito and Vázquez [5] establishes existence, uniqueness, and positivity properties of traveling waves and also analyzes the asymptotic behavior of solutions, proving that sufficiently rapidly decaying perturbations of the unstable state (say, 0 if $\theta<0$) converge to the stable state ($1$ if $\theta<0$) inside a ballistically growing region, with the rate of growth determined by the minimal wave speed. It would be interesting to extend such results to the present context.
@@ 1.6. Related Literature
The proof in [6] is based on the explicit construction of sub- and supersolutions of (11), which, by monotonicity, can be used to control solutions, or what is referred to as a barrier argument in the PDE literature. The same strategy was employed by Chen, Duquesne, and Shi in [13], who derived the same $\text{Beta}(2,2)$ limit theorem as in Corollary 1 above, albeit under slightly more restrictive assumptions on $(f_{+},f_{-})$.
@@ 2. Preliminaries
2.1. Extended-Real-Valued Random Variables
Note that the definition (9) still makes sense if the input random variables are permitted to take the values $\infty$ and -$\infty$. It will be convenient to extend the definition to include such extended-real-valued random variables. To that end, for an extended-real-valued random variable $X$, define the CDF $F:\mathbb{R}\to[0,1]$ via the rule
@@ 2.1. Extended-Real-Valued Random Variables
$CDF(\overline{\mathbb{R}})=\{F:\mathbb{R}\to[0,1]\,\,\text{increasing, right-continuous}\}.$
Using standard extended-real-number arithmetic, from now on, the formula (10) defining $T$ will be understood to hold in the domain $CDF(\overline{\mathbb{R}})$.
@@
It is useful to note that $T$ is continuous with respect to vague convergence. Recall that $\{F_{n}\}$ converges vaguely to $F$ if $F_{n}(x)\to F(x)$ at each point of continuity $x$ of $F$.
@@
It suffices to recall that this mode of convergence for $(F_{n})_{n\in\mathbb{N}}$ is equivalent to the existence of a probability space $\mathbb{P}$ supporting a sequence $\{(X_{1,n},X_{2,n})\}$ and a vector $(X_{1},X_{2})$ such that, for each $n$, $X_{1,n}$ and $X_{2,n}$ are i.i.d. with CDF $F_{n}$; $(X_{1},X_{2})$ are i.i.d. with CDF $F$; and $(X_{1,n},X_{2,n})\to(X_{1},X_{2})$ $\mathbb{P}$-almost surely. (Note that the possibility that $X_{1}$ or $X_{2}$ are infinite poses no difficulties here.) Letting $Y_{n}=\Phi(X_{1,n},X_{2,n})$ and $Y=\Phi(X_{1},X_{2})$, where $\Phi$ is sampled independently of $\{(X_{1,n},X_{2,n})\}$ and $(X_{1},X_{2})$, the boundedness and continuity of $f_{+}$ and $f_{-}$ are enough to deduce that $Y_{n}\to Y$ almost surely as $n\to\infty$. By definition of $T$, this implies $TF_{n}\to TF$. ∎
@@ 2.4. Monotonicity
$S$ is defined analogously to $T$: Given two CDF’s $F,G\in CDF(\overline{\mathbb{R}})$, let $X_{F}$ and $X_{G}$ be independent extended-real-valued random variables distributed according to $F$ and $G$, respectively, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{F},X_{G})$. Let $S(F,G)$ denote the CDF of the random variable $Y$ defined as follows:
@@
On the other hand, if $X^{-}_{1}\leq X_{2}\leq X^{+}_{1}$, then invoking first (13) and then the fact that $f_{+}$ is nonincreasing (see (12)), one finds
@@
Using coupling arguments very similar to the one employed above, one readily deduces that the map $T=T(p,\mathbf{P})$ is nondecreasing with respect to $p$ and $f_{+}$ and nonincreasing with respect to $f_{-}$. The result is stated next for the sake of precision.
@@
(i) For any $p_{1}\leq p_{2}$, $T(p_{1},\mathbf{P})F\leq T(p_{2},\mathbf{P})F$ for all $F\in CDF(\overline{\mathbb{R}})$.
(ii) If $\mathbf{P}$ and $\mathbf{\tilde{P}}$ are measures on $BC([0,\infty)\times BC([0,\infty)$ for which it is possible to find a coupling of the corresponding random variables $(f_{+},f_{-})$ and (f~+,f~)($\tilde$$_{+},\tilde{f}_{-}$) such that $f_{+}\leq\tilde{f}_{+}$ and $f_{-}\geq\tilde{f}_{-}$, then $T(p,\mathbf{P})F\leq T(p,\mathbf{\tilde{P}})F$ for each $F\in CDF(\overline{\mathbb{R}})$.
2.5. Absolute Continuity
@@ 3. Evolution of the CDF
This section concerns the evolution equation (11) satisfied by CDF’s. To this end, it will be useful to decompose the random variable $Y$ in the definition of $T$ (see (10)) in a specific way. Given $F\in CDF(\overline{\mathbb{R}})$, let $X_{1},X_{2}$ be i.i.d. extended-real-valued random variables distributed according to $F$, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{1},X_{2})$. Define the extended-real-valued random variable $Z$ by
@@
Using the explicit formulas (31) and (30), one obtains very concrete expressions for the action of $\mathscr{L}$ on smooth CDF’s, as stated in the next result.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is differentiable, then
@@ 3.1. Proof of Proposition 5
In [21, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph. Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the equation.
@@ 3.2. Proof of Proposition 6
Suppose that $F\in CDF(\overline{\mathbb{R}})$ is continuously differentiable. To obtain the formula (33) for $\mathscr{L}F$ in this case, it suffices to establish that
@@
In particular, in the analysis that follows, only the case where $Z=X_{2}$ needs to be treated. Further, there is no loss of generality assuming that $x\in\mathbb{R}$ as the integral vanishes for $x\in\{-\infty,+\infty\}$ because $F(x_{1}+g_{+}(+\infty-x_{1}))=F(+\infty)$.
@@ 3.2.1. Case: $\Theta=1$
$\{Z\leq x<Y,\,\,X_{1}<X_{2},\,\,\Theta=1\}$ $=\{-\infty<X_{1}<X_{2}\leq x,\,\,g_{+}(x-X_{1})<X_{2}-X_{1}\}$
$=\{-\infty<X_{1}<x,\,\,X_{1}+g_{+}(x-X_{1})<X_{2}\leq x\}.$
Therefore, since $f_{+}$ is independent of $(X_{1},X_{2})$ and $\mathbb{P}\{X_{1}\leq X_{2}\}=\mathbb{P}\{X_{1}<X_{2}\}$ by the continuity of $F$,
@@ 3.3. Proof of Proposition 7
$\lim_{x\to+\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}$ $=\lim_{x\to+\infty}\left(\mathbb{P}\{Z\leq x\,\mid\,\Theta=1\}-\mathbb{P}\{Y\leq x\,\mid\,\Theta=1\}\right)$
$=\mathbb{P}\{Z=+\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Y=+\infty\,\mid\,\Theta=1\}=0$
since $\{Y=+\infty\}=\{Z=+\infty\}$ by the boundedness of $f_{+}$ and $f_{-}$. By symmetry,
$\lim_{x\to-\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=0.$
@@
Next, the proof of (ii): Suppose that $F,G\in CDF(\overline{\mathbb{R}})$ differ by a constant, hence $G=F+c$ for some $c\in\mathbb{R}$. First, if $F$ is smooth, then equation (33) applied to $F$ and $F+c$ implies $\mathscr{L}(F+c)=\mathscr{L}F$. In general, there is a sequence $\{F_{n}\}$ of smooth CDF’s that converges vaguely to $F$. By Proposition 1, $\{TF_{n}\}$ and $\{T(F_{n}+c)\}$ converge vaguely to respective limits $TF$ and $T(F+c)$. At the same time, since the polynomial P(q)=q+(12p)q(1q)P(q)=q+(1-2p)q(1-q) is smooth, the functions $\{P(F_{n})\}$ converge vaguely to $\{P(F)\}$. Therefore, for a dense set of points $x\in\mathbb{R}$,
@@
and, similarly, $TF(-\infty)=P(q)$. Therefore, being nondecreasing, $TF$ must be a constant function, equal to $P(q)$. Identifying $F$ with $q$, this proves that $T$ maps $q$ to the constant $Tq=P(q)$, as claimed. ∎

### v3

@@ 1.2. General Setting
Without loss of generality, the pair $(f_{+},f_{-})$ will be assumed to be independent of $\Theta$ henceforth.
@@
(9) $X^{(n)}\overset{d}{=}\Phi(X^{(n-1)}_{1},X^{(n-1)}_{2})$,
where $X^{(n-1)}_{1},X^{(n-1)}_{2}$ are i.i.d. copies of $X^{(n-1)}$, which are independent of $\Phi$. Throughout the paper, equality is only required to hold in law; questions of couplings or constructions of $\{X^{(n)}\}$ will not be treated.
@@ 1.4.1. Main Results
(15) $\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<+\infty,$
@@
$\sigma=-\int_{0}^{+\infty}\mathbf{E}[|g_{+}(s)-s|]\,ds+\int_{0}^{+\infty}\mathbf{E}[|g_{-}(s)-s|]\,ds.$
@@
(16) $\int_{0}^{+\infty}(1+s)\mathbf{E}[|g_{+}(s)-s|+|g_{-}(s)-s|]\,ds<+\infty,$
let $a\geq 0$ be the constant defined by
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (16) holds, $\sigma=0$, and $a>0$. Fix a $\theta\in\mathbb{R}$ and let $p^{(N)}=1/2+\theta N^{-1}$. For each $N\in\mathbb{N}$, let $\{X^{(N,n)}\}$ be a solution of the RDE
@@ 1.4.2. Idea of the Proof
Consistency, on the other hand, refers to the behavior of the equation under rescaling. Specifically, Proposition 10 below shows that if $G$ is a smooth CDF, rescaled according to $G_{\delta}(x)=G(\delta x)$, then
(23) $(\mathscr{L}G_{\delta})(\delta^{-1}x)=\delta^{2}\sigma|\partial_{x}G(x)|^{2}+\delta^{3}a|\partial_{x}G(x)|\partial_{x}^{2}G(x)+\cdots\quad\text{as}\,\,\delta\downarrow 0.$
@@
which motivates the choices $\alpha=2$ and $\alpha=3$ used above and the appearance of the PDE’s (19) and (22). The approach of [7] provides a robust framework for making such a formal argument rigorous.
1.5. Variants of the Fisher-KPP Equation when $p\neq 1/2$
The results of this work suggest interesting new questions away from the critical point $p=1/2$. In view of the asymptotic expansion (23) and the PDE’s obtained in Theorems 1 and 2, the equation (11) shares some similarities with the reaction-diffusion equation
@@ 1.5. Variants of the Fisher-KPP Equation when $p\neq 1/2$
A number of works in the PDE literature considered the equation (24) in the case when $\sigma=0$ and $a>0$. Using an auxiliary ODE, Enguiça, Gavioli, and Sanchez [17] established the existence of traveling waves propagating at any large-enough speed, as in the classical Fisher-KPP equation. The work of Audrito and Vázquez [5] establishes existence, uniqueness, and positivity properties of traveling waves and also analyzes the asymptotic behavior of solutions, proving that sufficiently rapidly decaying perturbations of the unstable state (say, 0 if $\theta<0$) converge to the stable state ($1$ if $\theta<0$) inside a ballistically growing region, with the rate of growth determined by the minimal wave speed. It would be interesting to adapt the techniques of [5] to the study of (11) away from the critical point.
@@ 1.6. Related Literature
The proof in [6] is based on the explicit construction of sub- and supersolutions of (11), which, by monotonicity, can be used to control solutions, or what is referred to as a barrier argument in the PDE literature. The same strategy was employed by Chen, Duquesne, and Shi in [13], who independently derived the $\text{Beta}(2,2)$ limit theorem of Corollary 1 above, albeit under slightly more restrictive assumptions on $(f_{+},f_{-})$.
@@ 2. Preliminaries
2.1. Extended Real-Valued Random Variables
Note that the definition (9) still makes sense if the input random variables are permitted to take the values $\infty$ and -$\infty$. It will be convenient to extend the definition to include such extended real-valued random variables. To that end, for an extended real-valued random variable $X$, define the CDF $F:\mathbb{R}\to[0,1]$ via the rule
@@ 2.1. Extended Real-Valued Random Variables
$CDF(\overline{\mathbb{R}})=\{F:\mathbb{R}\to[0,1]\,\,\text{nondecreasing, right-continuous}\}.$
Using standard extended real number arithmetic, from now on, the formula (10) defining $T$ will be understood to hold in the domain $CDF(\overline{\mathbb{R}})$.
@@
It is useful to note that $T$ is continuous with respect to vague convergence. Recall that $\{F_{n}\}$ converges vaguely to $F$ if $F_{n}(x)\to F(x)$ at each point of continuity $x\in\mathbb{R}$ of $F$.
@@
It suffices to recall that this mode of convergence for $\{F_{n}\}$ is equivalent to the existence of a probability space $\mathbb{P}$ supporting a sequence $\{(X_{1,n},X_{2,n})\}$ and a vector $(X_{1},X_{2})$ such that, for each $n$, $X_{1,n}$ and $X_{2,n}$ are i.i.d. with CDF $F_{n}$; $(X_{1},X_{2})$ are i.i.d. with CDF $F$; and $(X_{1,n},X_{2,n})\to(X_{1},X_{2})$ $\mathbb{P}$-almost surely. (Note that the possibility that $X_{1}$ or $X_{2}$ are infinite poses no difficulties here.) Letting $Y_{n}=\Phi(X_{1,n},X_{2,n})$ and $Y=\Phi(X_{1},X_{2})$, where $\Phi$ is sampled independently of $\{(X_{1,n},X_{2,n})\}$ and $(X_{1},X_{2})$, the boundedness and continuity of $f_{+}$ and $f_{-}$ are enough to deduce that $Y_{n}\to Y$ almost surely as $n\to\infty$. By definition of $T$, this implies $TF_{n}\to TF$. ∎
@@ 2.4. Monotonicity
$S$ is defined analogously to $T$: Given two CDF’s $F,G\in CDF(\overline{\mathbb{R}})$, let $X_{F}$ and $X_{G}$ be independent extended real-valued random variables distributed according to $F$ and $G$, respectively, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{F},X_{G})$. Let $S(F,G)$ denote the CDF of the random variable $Y$ defined as follows:
@@
On the other hand, if $X^{-}_{1}\leq X_{2}\leq X^{+}_{1}$, then invoking first the fact that $f_{+}$ is nonincreasing (see (12)) and then (13), one finds
@@
Using coupling arguments very similar to the one employed above, one readily deduces that the map $T=T(p,\mathbf{P})$ is nonincreasing with respect to $p$ and $f_{+}$ and nondecreasing with respect to $f_{-}$. The result is stated next for the sake of precision.
@@
(i) For any $p_{1}\leq p_{2}$, $T(p_{1},\mathbf{P})F\geq T(p_{2},\mathbf{P})F$ for all $F\in CDF(\overline{\mathbb{R}})$.
(ii) If $\mathbf{P}$ and $\mathbf{\tilde{P}}$ are measures on $BC([0,\infty)\times BC([0,\infty)$ for which it is possible to find a coupling of the corresponding random variables $(f_{+},f_{-})$ and (f~+,f~)($\tilde$$_{+},\tilde{f}_{-}$) such that $f_{+}\leq\tilde{f}_{+}$ and $f_{-}\geq\tilde{f}_{-}$, then $T(p,\mathbf{P})F\geq T(p,\mathbf{\tilde{P}})F$ for each $F\in CDF(\overline{\mathbb{R}})$.
2.5. Continuity
@@ 3. Evolution of the CDF
This section concerns the evolution equation (11) satisfied by CDF’s. To this end, it will be useful to decompose the random variable $Y$ in the definition of $T$ (see (10)) in a specific way. Given $F\in CDF(\overline{\mathbb{R}})$, let $X_{1},X_{2}$ be i.i.d. extended real-valued random variables distributed according to $F$, and sample the data $(\Theta,f_{+},f_{-})$ independently of $(X_{1},X_{2})$. Define the extended real-valued random variable $Z$ by
@@
Using the explicit expressions (30) and (31), one obtains very concrete expressions for the action of $\mathscr{L}$ on smooth CDF’s, as stated in the next result.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is absolutely continuous, then
@@ 3.1. Proof of Proposition 5
In [21, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph. Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the evolution equation (11).
@@ 3.2. Proof of Proposition 6
Suppose that $F\in CDF(\overline{\mathbb{R}})$ is absolutely continuous and fix $x\in\mathbb{R}$. To obtain the formula (33) for $\mathscr{L}F(x)$ in this case, it suffices to establish that
@@
In particular, in the analysis that follows, only the case where $Z=X_{2}$ needs to be treated.
@@ 3.2.1. Case: $\Theta=1$
$\{Z\leq x<Y,\,\,X_{1}<X_{2},\,\,\Theta=1\}$
$\qquad=\{-\infty<X_{1}<X_{2}\leq x,\,\,g_{+}(x-X_{1})<X_{2}-X_{1},\,\,\Theta=1\}$
$\qquad=\{-\infty<X_{1}<x,\,\,X_{1}+g_{+}(x-X_{1})<X_{2}\leq x,\,\,\Theta=1\}.$
Therefore, since $(\Theta,f_{+},X_{1},X_{2})$ are mutually independent and $\mathbb{P}\{X_{1}\leq X_{2}\}=\mathbb{P}\{X_{1}<X_{2}\}$ by the continuity of $F$,
@@ 3.3. Proof of Proposition 7
$\lim_{x\to+\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}$ $=\lim_{x\to+\infty}\left(\mathbb{P}\{Y>x\,\mid\,\Theta=1\}-\mathbb{P}\{Z>x\,\mid\,\Theta=1\}\right)$
$=\mathbb{P}\{Y=+\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Z=+\infty\,\mid\,\Theta=1\}=0$
since $\{Y=+\infty\}=\{Z=+\infty\}$ by the boundedness of $f_{+}$ and $f_{-}$. Similarly,
$\lim_{x\to-\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}=\mathbb{P}\{Z=-\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Y=-\infty\,\mid\,\Theta=1\}=0$
as $\{Y=-\infty\}=\{Z=-\infty\}$. By symmetry,
$\lim_{x\to+\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=\lim_{x\to-\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=0.$
@@
Next, the proof of (ii): Suppose that $F,G\in CDF(\overline{\mathbb{R}})$ differ by a constant, hence $G=F+c$ for some $c\in\mathbb{R}$. First, if $F$ is smooth, then equation (33) applied to $F$ and $F+c$ implies $\mathscr{L}(F+c)=\mathscr{L}F$. In general, there is a sequence $\{F_{n}\}$ of smooth CDF’s that converges vaguely to $F$ and such that $F_{n}(\pm\infty)=F(\pm\infty)$ for any $n$. By Proposition 1, $\{TF_{n}\}$ and $\{T(F_{n}+c)\}$ converge vaguely to respective limits $TF$ and $T(F+c)$. At the same time, since the polynomial P(q)=q+(12p)q(1q)P(q)=q+(1-2p)q(1-q) is smooth, the functions $\{P(F_{n})\}$ converge vaguely to $\{P(F)\}$. Therefore, for a dense set of points $x\in\mathbb{R}$,
@@
and, similarly, $TF(-\infty)=P(q)$. Therefore, being nondecreasing, $TF$ must be a constant function, equal to $P(q)$. Identifying $F$ with $q$, this proves that $T$ maps $q$ to the constant $P(q)$, as claimed. ∎